Write the slope-intercept equation for the line that passes through (-7, 6) and is perpendicular to -7x + 9y = -2 Please show all of your work.
Answer by mananth(16946)

-7 x + 9 y = -2
Find the slope of this line

9 y = 7 x + 2
Divide by 9
y = 7/9 x 0.22
Compare this equation with y=mx+b
slope m = 0.78

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be -1.29
m= -1.29 ,point (-7,6)
Find b by plugging the values of m & the point in
y=mx+b
6=9+b
b=-3
m=-1.29
The required equation is y=-1.29 x-3